One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
This is where I started with trying to make an equation then I gave up on that and tried doing it the old fashioned way of using my thinking skills and I came up with 17 as my answer. Is it right?
x>12
39=y(2x)

Answer by checkley75(3666)

Y=X+12
X+2Y=39
X+2(X+12)=39
X+2X+24=39
3X=39-24
3X=15
X=15/3
X=5 ANSWER.
Y=5+12
Y=17 ANSWER.